How do u solve 0.3÷3
Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

How do u solve 0.3÷3
----
(3/10)/3 = (3/10)(1/3) = 1/10 = 0.1
======================================
Cheers,
Stan H.
============

Answer by solver91311(24713)



I open my calculator

I press the . (decimal point) key

I press the 3 key

I press the / (division) key

I press the 3 key

I press the = key

I read the answer in the display.

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism